OpenStudy (anonymous):

Which of the following are solutions to the equation cos^2 (2x) - 1/4 = 0? Check ALL that apply. A. 5 pi/ 6 B. 11pi / 3 C. pi / 6 D. 12pi / 6

OpenStudy (anonymous):

From doing some current calculations. I came to the conclusion of C. Not sure if it is right or if there are anymore answers

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm Cos}^2 (2x) - \frac{1}{4} = 0 }\) \(\large\color{black}{ \displaystyle {\rm Cos}^2 (2x)= \frac{1}{4} }\) \(\large\color{black}{ \displaystyle \left(~~ {\rm Cos} (2x)~\right)^2= \left(\frac{1}{2}\right)^2 }\)

OpenStudy (solomonzelman):

hope that helps...

OpenStudy (solomonzelman):

(take the square root of both sides, and don't forget the \(\pm\) )

OpenStudy (anonymous):

Ok! I will try to solve from there and I'll tell you what I get

OpenStudy (solomonzelman):

sure:)

OpenStudy (solomonzelman):

(oh, there is only 1 incorrect option)

OpenStudy (anonymous):

So \[\cos(2x) = \pm \frac{ 1 }{ 2 }\] If I remember how to do this properly, I think I find what 1/2 is on the unit circle, which would be 60? Then divide 60 by 2 and get 30 degrees?

OpenStudy (anonymous):

So there are 3 correct answers?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

30 degrees is one of them (as you said)

OpenStudy (anonymous):

Ok! Now I do not know where to do from here

OpenStudy (anonymous):

We already have \[\frac{ \pi }{ 6}\] as one answer

OpenStudy (solomonzelman):

you can plug the rest of the options, if you don't want to go ahead and generate the solution sets for -1/2 and +1/2.

OpenStudy (solomonzelman):

yes π/6 is correct

OpenStudy (anonymous):

I don't think I plugged the numbers in properly, but I got 12pi/6 and 5pi/6?

OpenStudy (solomonzelman):

12π/6 is not right

OpenStudy (solomonzelman):

5π/6 is right

OpenStudy (anonymous):

Ok! So how would I plug them in? I know I did it wrong haha

OpenStudy (solomonzelman):

(click alt, and hold it click 2 2 7 respectively on your number pad on the left if you have one release alt: you get π)

OpenStudy (solomonzelman):

you would just plug the answer choices instead of x, into that equation cos²(2x)-1/4=0

OpenStudy (anonymous):

Oh and what gets you = 0 right?

OpenStudy (anonymous):

\[2 \cos ^22x=\frac{ 1 }{ 2 }\] \[1+\cos 4x=\frac{ 1 }{ 2 }\] \[\cos 4x=\frac{ 1 }{ 2 }-1=-\frac{ 1 }{ 2 }=-\cos \frac{ \pi }{ 3 }=\cos \left(2n \pi+ \pi \pm \frac{ \pi }{ 3 } \right)\] where n is an integer. \[4x=\left( 2n+1 \right)\pi \pm \frac{ \pi }{ 3 }\] \[x=\frac{ 1 }{ 4 }\left\{ \left( 2n+1 \right)\pi \pm \frac{ \pi }{ 3 } \right\}\] now you can check by plugging n=0,1,2,....

OpenStudy (solomonzelman):

cos(4x) :) I haven't thought of that one....

OpenStudy (anonymous):

So the answers would be \[\frac{ 5\pi }{ 6}, \frac{ 11\pi }{ 3 }, and \frac{ \pi }{ 6 }?\]

OpenStudy (solomonzelman):

yes. use ~ for space

OpenStudy (solomonzelman):

yw

OpenStudy (anonymous):

Thank you for your time and the tips :)